A 20.0 L helium balloon was filled outdoor at a temperature of 5.0 oC. What temperature in oC is required to change the volume of the balloon to 31.6 L ? 

 Answer in 3 sig-dig number without unit oC



1
Expert's answer
2021-03-22T06:32:39-0400

V1/V2=T1/T2

V1 = 20 L

V2= 31.6

T1= 5.0

20/31.6= 5/X

0.6329114= 5/X

0.6329114X =5

X= 5/0.6329114

= 7.90